OpenStudy (anonymous):

4. Box A contains 8 red and 6 blue tokens. Box B contains 5 red and 11 blue tokens. One token is randomly taken from each box. a) What is the probability of getting a blue token from Box A and a red token from Box B? b) What is the probability of getting a blue token and a red token? c) What is the probability that both tokens are of the same colour?

OpenStudy (kropot72):

a) The probability of getting a blue token from box A is 6/14, and the probability of getting a red token from box B is 5/16. The two events are independent, therefore the probability of both occurring is: \[\frac{6\times5}{14\times16}=you\ can\ calculate\]

OpenStudy (anonymous):

83/112. How about part b?

OpenStudy (kropot72):

For part a), how did you get 83 for the numerator of your proposed answer?

OpenStudy (anonymous):

oops it's actually 15/112.

OpenStudy (kropot72):

Correct! b) The probability of getting a red token from box A and a blue token from box B is: \[\frac{8\times11}{14\times16}=you\ can\ calculate\] When you have found this probability, add it to the value of probability for part a). This is done because the two events, 'blue A / red B' and 'red A / blue B', are mutually exclusive.

OpenStudy (anonymous):

59/112? Thanks!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (kropot72):

c) You need to work out the probabilities of 'blue A / blue B' and 'red A / red B', and then add them.

OpenStudy (anonymous):

mind being more specific? I don't get it...

OpenStudy (anonymous):

is it 53/112?

OpenStudy (kropot72):

P(blue A) = 6/14 P(blue B) = 11/16 \[P(blue\ A /\ blue\ B)=\frac{6\times11}{14\times16}\] P(red A) = 8/14 P(red B) = 5/16 \[P(red\ A\ /\ red\ B)=\frac{8\times5}{14\times16}\] P('blue A / blue B' or 'red A / red B') is given by: \[(\frac{6\times11}{14\times16})+(\frac{8\times5}{14\times16})=you\ can\ calculate\]

OpenStudy (kropot72):

Your answer is correct!